OpenStudy (anonymous):

I NEED HELP with these type of problems. If you can show how you got the answer, it will help me a lot! Find the x and y intercepts of the line represented by the equation: y = -5x +5. x intercept = ( , 0 ) y intercept = ( 0, )

OpenStudy (jack1):

y = -5x +5 y intercept: so x = 0 y = -5x +5 y = -5*0 +5 y = 0 +5 y = 5 x intercept: so y = 0 y = -5x +5 0 = -5x +5 can u solve from here?

OpenStudy (jack1):

@buenotaco ?

OpenStudy (anonymous):

Yup, I got it. (1,0) (0,5).